OpenStudy (anonymous):

An electron in Li2+ falls from n = 2 to the ground state. Calculate the wavenumber, (νˉ), of the emitted photon.

OpenStudy (anonymous):

U knw Rydbergs Formula....Use it: \[\frac{ 1 }{ \lambda } = 109677z^2(\frac{ 1 }{ n1^2 }-\frac{ 1 }{ n2^2) }\] \[ (νˉ) = \frac{ 1 }{ \lambda }\]

OpenStudy (anonymous):

n1 = 1 n2 = 2 z = 3

OpenStudy (anonymous):

Hope U can Solve the Rest..)